OpenStudy (anonymous):

i put (170/ 360) * (3.14*9.1*9.1/4) and none of my answers match , does anyone know the problem????? Find the area of a sector with a central angle of 170° and a diameter of 9.1 cm. Round to the nearest tenth. A. 122.9 cm2 B. 30.7 cm2 C. 8.6 cm2 D. 3.4 cm2

jimthompson5910 (jim_thompson5910):

the radius is 9.1/2 = 4.55

jimthompson5910 (jim_thompson5910):

so you use that instead of the 9.1

OpenStudy (anonymous):

i got 1 after i multiplied it all and then divided :/ i had 11050/1440 was that correct?

jimthompson5910 (jim_thompson5910):

now divide the two

OpenStudy (anonymous):

i got 1

jimthompson5910 (jim_thompson5910):

that's not correct 11050/1440 is not 1

OpenStudy (anonymous):

sorry i pressed the wrong symbol , i got 7.673.....

OpenStudy (anonymous):

which would make it C , correct?

jimthompson5910 (jim_thompson5910):

no, one sec

jimthompson5910 (jim_thompson5910):

oh i see what's wrong, not sure why i didn't catch it til now A = (angle/360)*(pi*r*r) A = (170/360)*(3.14*4.55*4.55) A = (0.47222222)*(3.14*4.55*4.55) A = (0.47222222)*(65.00585) A = 30.697206799987 A = 30.7 So the answer is actually choice B

OpenStudy (anonymous):

thanks a million for your help! :)

jimthompson5910 (jim_thompson5910):

yw